Thank you as always, based on this problem i got another question

How did you get that LCD was 6.25? Was it by multiplying 1, 1.25, and 5?
based on the equation 1/x + 1/1.25x = 1/5
Please explain the process .. Thanks as always.

Answer by stanbon(75887)

1/x + 1/1.25x = 1/5
Right; the lcm = 5*1.25x
--------------------------
Multiply thru to get rid of the denominators:
(5*1.25) + 5 = 1.25x
11.25 = 1.25x
x = 9
====================
Cheers,
Stan H.